OpenStudy (anonymous):

Which equation is in standard form and represents a line with slope –1 through the point (–3, 5)?

OpenStudy (gorv):

y-y1=slope(x-x1) (x1,y1)=(-3,5)

OpenStudy (gorv):

y-5=-1(x+3) y-5=-x-3 y=-x+2

OpenStudy (gorv):

@blkahf